OpenStudy (anonymous):

Together, Larry and Mo require 4hrs 48min to pave a driveway. Alone, Larry would require 4hrs more than Mo. How long would it take Mo to do the job alone? Hint: convert minutes to

OpenStudy (goldphenoix):

Larry + Mo = 4 hours and 48 minutes. Larry = Mo + 4 hours. (M+4 hours) + M = 4 hours 48 minutes. (M+240minutes) + M = 288 minutes If you do the equation: M+240+M= 288 M+M= 48 2M=48 M=24 So Mo should take 24 minutes to do the job alone. Not 100% sure though.

OpenStudy (goldphenoix):

If you plug it in, Larry must do 4 hours or 240 minutes more than Mo. Since I'm assuming Mo takes 24 minutes to do the job alone, then Larry must take 264 minutes (240+24). Don't forget about Mo, so 264+24=284. Now i'm POSITIVE that Mo takes 24 minutes to do the job alone.